# Decoding AI Judgment: How LLMs Assess News Credibility and Bias

### v2

@@ Introduction
The protocol, detailed in the Methods section, unfolds as follows. An LLM agent first selects five out of six predefined evaluation criteria and ranks them by importance. Two additional agents then retrieve content: one downloads the homepage, while the other extracts up to two articles deemed informative for assessing credibility. The number of retrieved articles is not fixed, allowing the agent to autonomously select zero, one, or two. Subsequently, five additional agents each assess one of the selected criteria independently and in parallel. A final agent then aggregates these assessments into a structured summary and produces a binary reliability classification. This workflow enables a direct procedural comparison between LLMs and humans, both following the same evaluation steps in a controlled setting.
@@
To facilitate this comparison, we replicated the agentic protocol in a human subject experiment. We recruited $N=50$ participants and tasked them with assessing the reliability of news domains using the same structured procedure: selecting five criteria, ranking them, browsing the homepage and up to two articles, and finally producing a binary reliability judgment. Full details on participant recruitment and setup are provided in the Methods section.
@@
Figure 4 summarizes the results for the same set of 31 Italian-language news domains. Panel (a) compares the reliability ratings produced by LLMs and humans against NewsGuard’s classifications. LLMs, even when constrained to a more rigid evaluation pipeline and operating on non-English content, maintain consistency with earlier findings, though with lower accuracy. In contrast, human participants show no meaningful alignment with NewsGuard: reliable and unreliable domains are classified with roughly equal probability, suggesting that non-expert evaluators rely on different and less consistent heuristics. Panel (b) compares LLM and human ratings directly, treating human judgments as a referential baseline. While both groups agree on which outlets are unreliable, major divergence arises for reliable sources: LLMs classify as unreliable roughly 80% of the domains rated as reliable by humans.
@@
This asymmetry aligns with prior evidence that individuals are often more prone to reject accurate information than to believe falsehoods. A recent meta-analysis covering over 195,000 participants across 67 studies confirms this tendency—commonly referred to as skepticism bias—and shows that non-experts are especially likely to misclassify true information as false [45]. Our findings suggest that this bias persists even under structured and controlled evaluation.
@@
To explore the basis of this divergence, we examine the evaluation criteria selected by each group. Figure 4c shows the ranking distributions across the six available criteria (Table 1). Both groups consistently prioritize “Factual Accuracy”, defined as the extent to which the website’s content is “accurate and free from false or misleading information”. Nearly all participants—human and LLM—selected this criterion and ranked it first. However, this convergence masks deeper differences in reasoning processes. For human participants, prioritizing accuracy likely reflects deliberative reasoning based on content comprehension and analytical judgment [46]. In contrast, LLMs’ emphasis on accuracy emerges from learned statistical correlations rather than conceptual understanding.
@@
Beyond this shared top criterion, the hierarchies diverge sharply. LLMs consistently rank “Ownership Transparency” among the top three criteria, while humans rarely select it and often place it last. This is consistent with findings that most individuals—unlike professional fact-checkers—rarely engage in “lateral reading” to verify sources or consult external indicators of trustworthiness [47]. Even when constrained to internal browsing, LLMs exhibit a more structured verification behavior, possibly due to their exposure to factuality benchmarks during training. Human evaluators, by contrast, tend to rely on surface-level cues and intuitive judgments—especially when cognitive resources are limited [48, 49].
Human participants tend to prioritize rhetorical and stylistic cues, such as “Language Manipulation” and “Writing Professionalism”. Emotional tone is often interpreted as a sign of persuasive intent rather than factual reliability [24], and fluency of expression is linked to perceived truthfulness through the processing fluency heuristic [50]. LLMs, lacking affective or metacognitive grounding, treat these features as superficial markers, assigning them no epistemic weight.
Overall, this agentic comparison enables a more granular understanding of how LLMs operationalize credibility. What emerges is not just a difference in outcomes, but a structural divergence in evaluation construction. Humans blend intuitive and analytical processes [51]; LLMs rely on statistical reasoning, with no awareness, intentionality, or abstraction. Recognizing this distinction does not diminish the utility of LLMs in evaluative tasks, but it clarifies the boundaries of their competence. Delegating judgment to such systems may not replicate human reasoning—it may redefine it.
@@ Conclusions
This study examined how Large Language Models (LLMs) assess the reliability of news domains, comparing their evaluations to expert benchmarks and to evaluations produced by human participants following an identical protocol. While model outputs often align with expert classifications—especially when flagging unreliable sources—this apparent agreement conceals a deeper divergence in evaluative mechanisms.
@@
Our findings show that LLMs rely on lexical associations, statistical priors, and structural cues rather than contextual interpretation. This is particularly evident in the systematic asymmetry observed for reliable sources: right-leaning outlets are more frequently classified as unreliable, even in zero-content settings. When comparing LLM outputs to human judgments, we find further divergence: models tend to reject sources that humans rate as reliable, reflecting a pattern consistent with skepticism bias—an over-rejection of accurate information recently documented in large-scale studies.
@@
Despite both groups prioritizing “Factual Accuracy” as the most important criterion, the underlying mechanisms differ. Human participants likely interpret accuracy through content comprehension and pragmatic reasoning; LLMs, instead, derive it from statistical regularities encoded during training. This distinction becomes clearer in the secondary criteria: LLMs emphasize “Ownership Transparency”—aligning with professional fact-checking protocols—while humans give more weight to stylistic and rhetorical features such as tone and writing fluency. These preferences mirror known heuristics like processing fluency, whereby clarity and emotional neutrality enhance perceived truth.
@@
Such discrepancies underscore a structural difference between intuitive, context-aware human evaluation and the pattern-based, procedural logic of LLMs. As these systems are increasingly embedded in decision-making pipelines—moderation, classification, prioritization—it becomes critical to assess not just whether their outputs appear reasonable, but how their internal procedures operationalize normative categories like reliability and bias.
Our structured agentic framework enables such comparison by aligning inputs, tasks, and justification protocols across humans and LLMs. While our sample size limits generalizability, the controlled design and multilingual robustness provide a solid foundation for further research.
In sum, what appears as model “alignment” may reflect only superficial output convergence. Delegating judgment to LLMs risks embedding a new epistemic regime, one grounded not in deliberation, but in approximation.
Methods
@@
We follow a zero-shot, closed-book approach to prompting for all six LLMs we experiment with, by providing no example classifications and no definitions of “reliability”, respectively. Our prompt, which we use for all models, is split into two parts.
@@
When evaluating the LLMs’ ability to classify news outlets using only their domain names, we slightly altered the prompt in Fig. 5 by substituting the first paragraph with the text “Given the domain of a news outlet, you classify it as either ’reliable’ or ’unreliable’ based on its content.”, and by replacing all other occurrences of ‘HTML’ with ‘URL’.
@@
We implemented the agentic workflow for outlet reliability classification with Google’s Agent Development Kit (ADK) [54]. ADK is a Python toolkit that allows for developing and orchestrating agentic systems. In particular, for our study, we developed a multi-agent system where each sub-task in the reliability evaluation procedure is delegated to a dedicated agent. The implementation firstly relies on what ADK calls a “workflow agent”, specifically a “Sequential Agent”. This is not an LLM-powered agent, but rather a component in ADK that allows for orchestrating the other agents in a deterministic manner. Specific to our case, this agent is what enables our procedure to follow a well-defined path, as agents are called one after the other. We provide all specific prompts used in this workflow in the Supplementary Information.
The first agent is a “code agent”, that is, an LLM agent that performs actions by writing code and calling tools written in a programming language [55], Python in our case. Specifically, this agent is tasked with selecting five criteria to evaluate from a list of six, and to rank them from “most to least important for assessing reliability”. To avoid our chosen order of criteria influencing the model’s decision, we don’t include the list directly in the prompt. Instead, the agent calls a Python function that returns the criteria in a randomized order.
@@
Once the criteria are selected and the data collected, the workflow activates all agents tasked with evaluating the criteria, assigning one criterion per agent. These LLM agents are not provided with any coding tool, meaning they function analogously to any other LLM assistant when given a prompt. Further, they act in parallel and produce their assessment regardless of each other’s output. Each agent is asked to output a rating from 1 to 5, with higher scores corresponding to higher reliability, a written summary explanation with examples and quotes from the analyzed content, and a binary flag indicating whether the agent also analyzed the downloaded articles for evaluating the criterion. Additionally, we instruct a separate agent to analyze the news outlet’s political orientation on a 5-point scale from Left to Right.
@@
A total of 50 participants (28 females, 22 males; $\mu_{\text{age}}=28.4$, $\sigma_{\text{age}}=9.6$) took part in the in-person experiment sessions at the Department of Psychology of Sapienza University of Rome. All participants were recruited online through advertisements distributed via social media platforms (e.g., LinkedIn, Facebook) and by snowball sampling. Eligibility was limited to native Italian-speaking adults (aged $\geq 18$) with normal or corrected vision.
@@
Acknowledgements SERICS (PE00000014) under the NRRP MUR program funded by the European Union - NextGenerationEU, project CRESP from the Italian Ministry of Health under the program CCM 2022, PON project “Ricerca e Innovazione” 2014-2020, and PRIN Project MUSMA for Italian Ministry of University and Research (MUR) through the PRIN 2022.
Supplementary Information
Reliability and Political Orientation
In this section, we provide further benchmarks for reliability and political orientation classifications. Figure S1 reports the reliability classification performance against Media Bias/Fact Check’s (MBFC) “Credibility Rating”. The results are consistent with NewsGuard (Fig. 1a in the main text). Figure S2 shows how models classify political orientation against NewsGuard’s, in panel (a), and MBFC’s, in panel (b), ratings. All six LLMs show strong agreement with both, with only minor differences attributable to the varying granularity of political labels: a 5-point scale for the models, 3-point for NewsGuard, and 7-point for MBFC.
Figure S1: LLMs’ classification against MBFC. All models show strong agreement with MBFC’s Credibility Rating. However, some differences arise for sources with Medium credibility.
Figure S2: LLMs’ classification of political orientation. (a) Comparison against Newsguard; (b) Comparison against Media Bias/Fact Check (MBFC).
Keyword Analysis
In this section, we supplement the results of our keyword analysis with the three remaining models: DeepSeek, Llama 3.1, and Mistral. Figure S3 shows the rank-frequency distribution of “classification” and “determinant keywords” across models and reliability ratings. Figure S4 displays keywords used for both reliable and unreliable domains and compares their rank across the two classifications.
Figure S3: Rank-frequency distributions of keywords used by each LLM to describe domains. This figure supplements Fig. 2 in the main text by presenting the results for the three remaining models: DeepSeek, Llama 3.1, and Mistral.
Figure S4: Keywords’ rank among “reliable” and “unreliable” domains. This figure supplements Fig. 3 in the main text by presenting the results for the three remaining models: DeepSeek, Llama 3.1, and Mistral.
Agentic Framework
In this section, we provide all prompts employed in our agentic framework:
  • The prompt in Fig. S5 is used for the agent tasked with selecting the criteria to evaluate.
  • The prompts in Fig. S6 and Fig. S7 are used for the two agents instructed to scrape the homepage and up to two articles, respectively.
  • The prompt in Fig. S8 is the generic template used for all agents tasked with evaluating each reliability criterion. Table S1 lists the substitutions specific to each agent.
  • The prompt in Fig. S9 is used for the agent tasked with evaluating the news outlet’s political orientation.
You are an AI assistant responsible for selecting reliability evaluation criteria for news outlets. You must use the ‘get_criteria_list’ tool, which provides a randomized list of questions. Rank the list of questions in order from most to least important for assessing reliability. Then, select the first 5 questions. Output *only* the questions resulting from this ranking and final selection. The output must have this JSON format:
  "criteria": [
    "<most important question>",
    "<next most important question>",
    "... (5 total)"
Provide no explanations or comments. Just the questions.
Figure S5: Prompt used for the agent responsible for selecting the evaluation criteria.
You are an AI that evaluates a news outlet’s reliability. The user will provide the news outlet’s homepage in a JSON format like {"url": "webpage_url"} You are able to access the page using a specific tool. You must: 1. extract the URL of the homepage 2. use the ‘scrape_webpage’ tool to retrieve the contents of the webpage 4. output *only* a JSON with this schema:
    {"webpage": "contents of the webpage"}
Figure S6: Prompt used for the agent responsible for scraping the news outlet’s homepage.
You are an AI that evaluates a news outlet’s reliability. You can access the news outlet’s homepage under session state key ‘news_outlet_homepage’. Your task is to select 0, 1, or 2 articles that may be indicative of the news outlet’s reliability. You must: 1. analyze the contents of the homepage, which is provided in Markdown format 2. detect any news article and retrieve their URL (also add the domain if the URL does not include it) 3. scrape their contents with the ‘scrape_webpage’ tool Output *only* a JSON with this schema:
  "article_1": "contents of the first article,
                if you selected one article",
  "article_2": "contents of the second article,
                if you selected two articles"
or with an empty JSON if you selected no articles.
Figure S7: Prompt used for the agent responsible for analyzing the scraped homepage and selecting up to two articles to collect.
You are an AI that evaluates a news outlet’s reliability. You can retrieve the contents of the news outlet’s webpages from the ‘news_outlet_homepage’ and ‘news_outlet_articles’ session state keys. You must evaluate the question: “{question} Output *only* a JSON with keys: - ‘criterion_question - ‘also_analyzed_articles’: (boolean) whether, in addition to the homepage, you have also assessed specific articles for this evaluation - ‘rating’: from 1 ({low rating description}) to 5 ({high rating description}) - ‘summary’: your explanations with precise examples and quotes from the content you analyzed
Figure S8: Prompt template used for all agents responsible for evaluating the reliability criteria. Questions and rating descriptions specific to each agent are reported in Table S1.
You are an AI that evaluates a news outlet’s political orientation. You can retrieve the contents of the news outlet’s webpages from the ‘news_outlet_homepage’ and ‘news_outlet_articles’ session state keys. Output *only* a JSON with keys: - ‘rating’: one between ’left’, ’center-left’, ’center’, ’center-right’, ’right’ - ‘also_analyzed_articles’: (boolean) whether, in addition to the homepage, you have also assessed specific articles for this evaluation - ‘summary’: your explanations with precise examples and quotes from the content you analyzed
Figure S9: Prompt template used for all agents responsible for evaluating the reliability criteria. Questions and rating descriptions specific to each agent are reported in Table S1.
Criterion Question Rating = 1 Rating = 5
Author Accountability To what extent does the site provide the names of content authors, along with their biographies or contact information? no author information all author information
Factual Accuracy To what extent do you believe the content presented on the site is accurate and free from false or misleading information? false and misleading accurate
Language Manipulation To what extent does the site use emotionally charged, exaggerated, or manipulative language? manipulative and exaggerated language neutral language
Ownership Transparency To what extent does the site clearly declare who owns it and who provides funding for it? not transparent transparent
Perspective Diversity To what extent does the site present content offering diverse perspectives without ideological or political bias? bias and no diversity neutral and diverse perspectives
Writing Professionalism To what extent does the site adhere to grammatical rules and use a clear, consistent, and professional writing style? unprofessional or improper style clear and professional style
Table S1: Evaluation criteria for news outlets and corresponding replacements for the prompt in Fig. S8.

### v3

@@ Introduction
The protocol, detailed in the Methods section, unfolds as follows. First, an LLM agent selects five out of six predefined evaluation criteria and ranks them by importance. This selection happens before the LLM is exposed to any content from the news outlet apart from its URL, which is necessary to start the pipeline. However, further experiments we conducted, presented in Supplementary Information, suggest that exposure to the URL does not appear to influence this selection. Additionally, the criteria are provided to the model in randomized order. Two additional agents then retrieve content: one downloads the homepage, while the other extracts up to two articles deemed informative for assessing credibility, using the URLs of articles found in the homepage. This content is collected using a tool available to agents that downloads webpages in a well-structured Markdown document. The number of retrieved articles is not fixed, allowing the agent to autonomously select zero, one, or two. Subsequently, five additional agents independently evaluate the selected criteria, each using only the materials retrieved in the prior step. Their outputs include both a numeric reliability rating from 1 to 5 and a written assessment. A final agent then aggregates these assessments into a structured summary and produces a binary reliability classification, based upon the other agents’ evaluations. This workflow enables a direct procedural comparison between LLMs and humans, both following the same evaluation steps in a controlled setting.
@@
To facilitate this comparison, we replicated the agentic protocol in a human subject experiment. We recruited $N=50$ participants and tasked them with assessing the reliability of online news outlets using the same structured procedure: selecting five criteria, ranking them, browsing the homepage and up to two articles, and finally producing a binary reliability judgment. Full details on participant recruitment and setup are provided in Materials and Methods.
@@
Although a total of 37 Italian-language news outlet domains were evaluated by the recruited human participants, only 27 of these could also be evaluated using our agentic workflow, as reported in Supplementary Table S2. This limitation was due to webpage retrieval requests being blocked by some domains. Figure 4 summarizes the results for this subset of domains, while in Supplementary Fig. S7 we report the results from human evaluators for the full set. Panel (a) compares the reliability ratings produced by LLMs and humans against NewsGuard’s classifications. LLMs, even when constrained to a more rigid evaluation pipeline and operating on non-English content, maintain consistency with earlier findings, though with lower accuracy. In contrast, human participants show no meaningful alignment with NewsGuard: reliable and unreliable domains are classified with roughly equal probability, suggesting that non-expert evaluators rely on different and less consistent indicators. Panel (b) compares LLM and human ratings directly, treating human judgments as a referential baseline. While both groups agree on which outlets are unreliable, major divergence arises for reliable sources: LLMs classify as unreliable almost 80% of the domains rated as reliable by humans.
@@
This asymmetry aligns with prior evidence that individuals are often more prone to reject accurate information than to believe falsehoods. A recent meta-analysis covering over 195,000 participants across 67 studies confirms this tendency—commonly referred to as skepticism bias—and shows that non-experts are especially likely to misclassify true information as false [48]. Our findings suggest that this bias persists even under structured and controlled evaluation. By directly integrating our experiment with this literature, we show that LLMs reproduce some behavioral regularities observed in humans (e.g., skepticism bias) while diverging sharply in the indicators that guide their judgments.
@@
To explore the basis of this divergence, we examine the evaluation criteria selected by each group. Figure 4c shows the ranking distributions across the six available criteria (Table 1). Both groups consistently prioritize “Factual Accuracy”, defined as the extent to which the website’s content is “accurate and free from false or misleading information”. Nearly all participants—human and LLM—selected this criterion and ranked it first. However, this convergence masks deeper differences in reasoning processes. For human participants, prioritizing accuracy likely reflects deliberative reasoning based on content comprehension and analytical judgment [49]. In contrast, for LLMs this emphasis appears to be operationalized via lexical associations and patterns learned during training, as suggested by proxy analyses.
@@
Beyond this shared top criterion, the hierarchies diverge sharply. LLMs consistently rank “Ownership Transparency” among the top three criteria, while humans rarely select it and often place it last. This is consistent with findings that most individuals—unlike professional fact-checkers—rarely engage in “lateral reading” to verify sources or consult external indicators of trustworthiness [50]. LLM outputs display patterns consistent with more structured verification behavior, possibly reflecting exposure to factuality benchmarks during training. Human evaluators, by contrast, often rely on surface-level cues and intuitive judgments—especially when cognitive resources are limited[51, 52].
Compared to LLMs, human participants tend to prioritize rhetorical and stylistic cues, such as “Language Manipulation” and “Writing Professionalism”. Emotional tone is often interpreted as a sign of persuasive intent rather than factual reliability [26], and fluency of expression is linked to perceived truthfulness through the processing fluency heuristic [53] and the well-documented illusory truth effect, the tendency to perceive repeated information as more likely to be true regardless of its actual accuracy [54]. Given that LLMs lack affective or metacognitive grounding, these features appear to carry little epistemic weight in their evaluations, functioning mainly as superficial markers. Taken together, this integration with prior literature suggests that LLMs approximate some dimensions of human evaluation while omitting others. Overall, this agentic comparison enables a more granular understanding of how LLMs appear to operationalize credibility. What emerges is not just a difference in outcomes but a difference in the observable patterns through which evaluation processes are instantiated.
Humans blend intuitive and analytical processes [55]; LLM outputs, by contrast, reflect different patterns of evaluation shaped by statistical learning. Recognizing this distinction does not diminish the utility of LLMs in evaluative tasks, but it clarifies the boundaries of their competence. Our findings point to a shift from content-based deliberation toward plausibility-driven approximations when judgment is delegated to automated systems—a dynamic we term epistemia, where statistical plausibility risks replacing deliberative reasoning with the illusion of knowledge rather than its verification. Future research could link behavioral experiments, cognitive models, and automated evaluation pipelines to map where human and machine heuristics converge and where they fundamentally diverge.
@@ Conclusions
This study examined how Large Language Models (LLMs) operationalize core evaluative concepts—such as reliability and bias—when tasked with assessing online news outlets, comparing their judgments to expert benchmarks and to human participants following the same protocol. While model outputs often align with expert classifications—especially when flagging unreliable sources—this apparent agreement conceals a deeper divergence in the evaluative mechanisms themselves, raising broader questions about delegating judgment to automated systems in information environments already shaped by infodemics and platform-driven filtering.
@@
Consistent with our findings, LLMs operate through lexical associations, statistical priors, and structural cues, rather than genuine contextual interpretation. This statistical approximation produces systematic asymmetries: right-leaning outlets are disproportionately classified as unreliable. When comparing LLM outputs to human judgments, we find further divergence: models display a pattern reminiscent of the skepticism bias observed in humans—an over-rejection of accurate information documented in large-scale studies [56].
@@
By integrating our human experiment with existing work on credibility assessment and cognitive biases [26, 25, 48], we show that LLMs partially replicate behavioral regularities identified in psychology while relying on fundamentally different evaluative mechanisms. This link clarifies that what emerges is not simply an accuracy gap, but a structural shift in how evaluation itself is operationalized when judgment is delegated to automated systems. Both groups prioritize “Factual Accuracy” as the most important criterion, but the underlying mechanisms differ. Human participants likely interpret accuracy through content comprehension and pragmatic reasoning; LLMs, instead, derive it from statistical regularities encoded during training. This distinction becomes clearer in the secondary criteria: LLMs emphasize “Ownership Transparency”—aligning with professional fact-checking protocols—while humans give more weight to stylistic and rhetorical features such as tone and writing fluency. These preferences mirror known heuristics like processing fluency, whereby clarity and emotional neutrality enhance perceived truth.
@@
Such discrepancies underscore a structural difference between intuitive, context-aware human evaluation and the pattern-based, procedural mechanisms of LLMs. As these systems are increasingly embedded in decision-making pipelines—moderation, classification, prioritization—it becomes critical to assess not just whether their outputs appear reasonable, but how their internal procedures operationalize normative categories like reliability and bias. This is especially urgent in an information ecosystem already marked by infodemics, where the oversupply of low-quality or contradictory information erodes trust and amplifies polarization [18]. In this context, the rise of what we term epistemia—the illusion of knowledge emerging when plausibility replaces verification—illustrates the risk that statistical approximation could displace deliberative reasoning if adopted uncritically.
Whereas the first wave of research on social media emphasized the volume and velocity of information flows [17, 23], our findings highlight a second phase in which the problem shifts from information overload to the nature of judgment itself, as automated pipelines introduce epistemic opacity into evaluative processes. Our structured agentic framework enables such comparison by aligning inputs, tasks, and justification protocols across humans and LLMs. While our sample size limits generalizability, the controlled design and multilingual robustness provide a solid foundation for further research. Future work should examine how this epistemic shift interacts with governance, transparency, and human oversight, especially as automated judgment pipelines expand beyond content moderation into law, policy, and scientific evaluation.
In sum, the apparent alignment between LLMs and expert judgments may mask only superficial output convergence. Delegating evaluative tasks to these systems risks embedding frameworks driven by lexical and statistical associations rather than deliberative reasoning, amplifying existing information pathologies. Addressing this shift requires transparency, human oversight, and potentially new training paradigms that explicitly disentangle factual reliability from ideological or stylistic cues. Hybrid approaches that combine statistical models with explicit reasoning criteria or retrieval-based evidence may offer a promising way forward.
Materials and Methods
@@
We follow a zero-shot, closed-book approach to prompting for all six LLMs we experiment with, by providing no example classifications and no definitions of “reliability”, respectively. Additionally, for all models, we use the system prompt “You are a service that gives a reliability rating to news outlets.” and set the temperature parameter to 0. Our prompt, which we use for all models, is split into two parts.
@@
If the output political orientation label fell outside the specified 5-point scale, we reassigned it to “Center”. For all models, this affected none or fewer than 1% of the domains, except for Mistral Large 2, where approximately 4% of the domains received an out-of-scale label.
When evaluating the LLMs’ ability to classify news outlets using only their domain names, we slightly altered the prompt in Fig. 5 by substituting the first paragraph with the text “Given the domain of a news outlet, you classify it as either ‘reliable’ or ‘unreliable’ based on its content.”, and by replacing all other occurrences of ‘HTML’ with ‘URL’.
@@
We implemented the agentic workflow for outlet reliability classification with Google’s Agent Development Kit (ADK) [59]. ADK is a Python toolkit that allows for developing and orchestrating agentic systems. In particular, for our study, we developed a multi-agent system where each sub-task in the reliability evaluation procedure is delegated to a dedicated agent. The implementation firstly relies on what ADK calls a “workflow agent”, specifically a “Sequential Agent”. This is not an LLM-powered agent, but rather a component in ADK that allows for orchestrating the other agents in a deterministic manner. Specific to our case, this agent is what enables our procedure to follow a well-defined path, as agents are called one after the other. The initial prompt used to start the workflow is “Select the most appropriate reliability criteria among those provided and evaluate the reliability of {URL}. We provide all additional prompts used in this workflow in the Supplementary Information.
The first agent is a “code agent”, that is, an LLM agent that performs actions by writing code and calling tools written in a programming language [60], Python in our case. Specifically, it is a tool-augmented agent that is tasked with selecting five criteria to evaluate from a list of six, and to rank them from “most to least important for assessing reliability”. To avoid our chosen order of criteria influencing the model’s decision, we do not include the list directly in the prompt. Instead, the agent calls a Python function that returns the criteria in a randomized order.
@@
Once the criteria are selected and the data collected, the workflow activates all agents tasked with evaluating the criteria, assigning one criterion per agent. These LLM agents are not provided with any coding tool, meaning they function analogously to any other LLM assistant when given a prompt. Further, they produce their assessment independently from each other. Each agent is asked to output a rating from 1 to 5, with higher scores corresponding to higher reliability, a written summary explanation with examples and quotes from the analyzed content, and a binary flag indicating whether the agent also analyzed the downloaded articles for evaluating the criterion. Additionally, we instruct a separate agent to analyze the news outlet’s political orientation on a 5-point scale from Left to Right.
@@
We set the temperature to 0 for all LLM agents, except for the one responsible for selecting the evaluation criteria. For this agent, we instead used a temperature of 1 to introduce more variability in the outputs. However, as shown in Supplementary Information Figures S5 and S6, the prioritization of criteria observed in Fig. 4c remains largely consistent across the full range of temperatures available for Gemini 2.0 Flash, likely because the task is highly constrained rather than open-ended.
@@
A total of 50 participants (28 females, 22 males; $\mu_{\text{age}}=28.4$, $\sigma_{\text{age}}=9.6$) took part in the in-person experiment sessions at the Department of Psychology of Sapienza University of Rome. All participants were recruited online through advertisements distributed via social media platforms (e.g., LinkedIn, Facebook) and by snowball sampling. Eligibility was limited to native Italian-speaking adults (aged $\geq 18$) with normal or corrected vision. All human participants were non-experts with no prior training in credibility assessment, ensuring that judgments reflected naïve, uncoached heuristics rather than professional fact-checking protocols.
@@
Acknowledgements We really thank Luciano Floridi, Mattia Samory, Giovanni Pezzulo and the Hypnotoad for precious insights during the development of the work. The work is supported by IRIS Infodemic Coalition (UK government, grant no. SCH-00001-3391), SERICS (PE00000014) under the NRRP MUR program funded by the European Union - NextGenerationEU, project CRESP from the Italian Ministry of Health under the program CCM 2022, PON project “Ricerca e Innovazione” 2014-2020, and PRIN. This work was supported by the PRIN 2022 “MUSMA” - CUP G53D23002930006 - Funded by EU - Next-Generation EU – M4 C2 I1.1.
[…]
Ethics Statement The study protocol was approved by the Ethics Committee for Transdisciplinary Research of Sapienza University of Rome (Protocol ID: 339/2025).
[…]
Author contribution W.Q. designed research; E.L., J.N., and N.D.M. performed data extraction and analysis; A.S. and R.A. designed and conducted the human experiments; M.C., V.C., C.R.-A., and W.Q. supervised the work; and all authors contributed to writing the manuscript and provided critical feedback.